OpenStudy (anonymous):

Find the sum of the arithmetic series: 3 + 5 + 7 + ... + 227

OpenStudy (calculusxy):

What does the "..." mean?

OpenStudy (anonymous):

that there are other #'s in there

OpenStudy (calculusxy):

Ok trying to solve it

OpenStudy (anonymous):

ok :)

OpenStudy (calculusxy):

Ok I got 12, 995

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

a=3,d=5-3=2 tn=a+(n-1)d 227=3+(n-1)*2 227-3=2(n-1) 224/2=n-1 112+1=n n=113 \[Sn=\frac{ n \left( a+l \right) }{ 2 }=\frac{ 113\left( 3+227 \right) }{2 }\] solve and get the solution.

OpenStudy (anonymous):

I guess I got 12,995 too

OpenStudy (calculusxy):

Ok. So I did this equation: 3 + 5 + 7 + ... + 227 2x1+1 + 2x2+1. + 2x3+1 +....+ 113x2+1 1x113+2(1+2+3...113) 113+2x(1/2)x(113x114) 113+113x114

OpenStudy (anonymous):

u get same answer

OpenStudy (calculusxy):

continued... 113x(1+114) 113x115= 12,995

OpenStudy (calculusxy):

@music101 Yes I got the same answer.